OpenStudy (cheska_p):

The curved section of a horizontal highway is a circular unbanked arc of radius 740 m. If the coefficient of static friction between this roadway and typical tires is 0.40, what would be the maximum safe driving speed for this horizontal curved section of highway? Can someone please show me how to do this?

OpenStudy (danjs):

when going around the circle , the car would accelerate radially towards the center of that circle

OpenStudy (danjs):

, the car would have a net force then [F=m*a] in the horizontal direction But the car does not want to slide off the road, so have to balance that with the friction force between the tire material and road material

OpenStudy (cheska_p):

Yes, but how do you calculate the velocity without mass?

OpenStudy (danjs):

The radial acceleration can be a function of the velocity

OpenStudy (cheska_p):

I'm sorry I don't understand

OpenStudy (danjs):

mass can be found in the other dimension if you need it, the normal force is to weight

OpenStudy (cheska_p):

And how would you calculate normal force?

OpenStudy (danjs):

The acceleration towards the circle creates a Force, centripital acceleration = v^2 / r F = m*a = m*v^2/r

OpenStudy (astrophysics):

Maybe a FBD will help,  maybe something like this, sorry for the bad drawing but notice that friction and centripetal acceleration are towards the center of the circle